A travelling wave is represented by the function:

y(x,t) = 0.009 m sin (1.2m^-1x – 5.0s^-1t)

Find the following:

a. Amplitude

b. Wave number

c. Wavelength

d. Angular frequency

e. Frequency

f. wave speed


1
Expert's answer
2021-02-11T10:40:15-0500

a)

"A= 0.009\ m."

b)

"k=1.2\ m^{-1}."

c)

"\lambda=\dfrac{2\pi}{k}=\dfrac{2\pi}{1.2\ m^{-1}}=5.23\ m."

d)

"\omega=5.0\ \dfrac{rad}{s}."

e)

"f=\dfrac{\omega}{2\pi}=\dfrac{5.0\ \dfrac{rad}{s}}{2\pi}=0.79\ Hz."

f)

"v=f\lambda= 0.79\ Hz\cdot5.23\ m=4.13\ \dfrac{m}{s}."